# Noé Arnold, Juan Beccuti, Thunj Chantramonklasri, Matthias Hafner, Nicolas Oderbolz Center for Cryptoeconomics

### v1

@@ V Discussion
It is important to note that the modeling approach above does not consider entry or exit in the sense that we assume a fixed number of agents per ETH holder type. Further, the model does not allow all agents to switch between different staking methods. Based on these assumptions, the model framework may be able to model the short-run dynamics of the Ethereum staking market, but may fall short when considering the long-run perspective. Consequently, the results presented here reflect a short-term perspective on how the staking market might respond to changes in the issuance schedule. In the long run, the relative profitability of different staking solutions is likely to influence market shares, as new entrants or existing market participants opt for profitable staking methods and existing stakers exit unprofitable ones. In the following, we therefore frame the discussion of the above results within a short-term perspective.
V-A Short-term Perspective
@@
The results of the calibrated game theoretic model, where solo stakers are modeled as experiencing disutility from the variability of MEV revenues and where a fraction of individuals staking through centralized providers are assumed to be inattentive, are broadly consistent with the empirical evidence suggesting that solo stakers have a higher yield elasticity of staking supply. Under these assumptions, the staking supply of solo stakers is more responsive to changes in the issuance schedule than that of other types of stakers.
This result underscores the fact that examining cost structures alone is not enough to explain the behavior of staking market participants. Even with cost structures that would otherwise make solo stakers inelastic to changes in consensus yields, the competitive environment in the broader staking market can still make solo stakers vulnerable to being crowded out of the staking market when consensus yields fall. This is because even in a segregated staking market where existing solo stakers do not switch to other staking solutions at all, other market participants may be relatively insensitive to changes in consensus issuance and thus put pressure on solo stakers to reduce their staking supply when consensus yields fall. This result also has important implications for the broader economic policy mix of the Ethereum protocol. It suggests that policies that help improve the competitive standing of solo stakers relative to other staking methods, for example the implementation of MEV burn, could help mitigate the otherwise negative effects of an issuance reduction on the market share of solo stakers.
@@
Assuming that tee model specification with all extensions most closely reflects empirical reality, the market share of solo stakers is expected to decrease under the proposed issuance schedule, while the market share of centralized exchanges would increase. This suggests that a reduction in issuance could help limit the market share of liquid staking providers and the resulting dominance of LSTs within the Ethereum ecosystem, at least in the short run. At the same time, a reduction in issuance may come at the cost of increased market share for centralized providers. Combined with a declining market share of solo stakers, this could contribute to an overall higher degree of centralization in the Ethereum validator set.
V-B Long-term Perspective
@@
However, exits and market participants switching between staking methods as a result of diminished profitability are not the only factors that could lead to diminished solo staking and an increasing dominance of liquid staking solutions in the market. This is because, over time, the divergence in profits between the different staking methods leads to a redistributive effect between the different staking methods. If stakers are able to reinvest their profits from staking, and staking profits are comparatively low for solo staking, their market share will continue to diminish. To illustrate this effect, we use the Nash equilibria of the baseline model outlined in Table I as a starting point and chart the cumulative changes in staking levels and staking ratios over time, assuming stakers reinvest their profits (revenues minus costs) and do not reevaluate their optimal level of staking supply over time. The results are shown in Figure 3. Figure 3 depicts the cumulative percentage change in the total staking level relative to the starting point. Similarly, Figure 4 depicts the cumulative change in the share of total ETH supply being staked. As such, Figure 4 also takes into account the effects of dilution.
@@
Figure 3: Cumul. long-run effects of profit reinvestment (staking level)
Figure 4: Cumulative long-run effects of profit reinvestment (staking ratio)
VI Conclusion
We show that solo stakers are more responsive to changes in consensus issuance yields than ETH holders staking through centralized exchanges or liquid staking providers. As issuance declines, solo staking is expected to lose market share, while centralized exchanges gain dominance.
This heightened sensitivity of solo stakers is not solely explained by their cost structures but is instead driven by the competitive dynamics of the staking market. Other staking methods benefit from superior MEV access and DeFi yields, making them less responsive to issuance changes and creating a crowding-out effect on solo stakers. These findings underscore the importance of mechanisms such as MEV burn in mitigating the negative impact of issuance reductions on solo stakers.
@@ VI Conclusion
The present analysis has certain limitations, like the assumptions used in calibrating the model. While these assumptions align with existing literature, they remain open to debate. Future research should prioritize refining the estimation of cost structures across different staking methods, which could enhance the accuracy and robustness of the model framework.
@@
Still, we expect that the model developed here serves as a foundation for further research and extensions, providing a flexible framework for analyzing staking dynamics under various protocol adjustments. For tractability and compatibility with the broader model, we introduced several simplifying assumptions in the model extension with endogenous dSSP fees. However, we believe that our model could be expanded to facilitate a deeper exploration of the competitive dynamics between staking intermediaries.

### v2

@@ 6 Discussion
There are several limitations to the theoretical model developed in this paper. First, it is important to note that the model does not consider entry or exits of staking agents, i.e., we assume a fixed number of agents per ETH holder type. Further, the model does not allow stakers to switch between different staking methods. Based on these assumptions, the model framework may be able to model the short-run equilibrium dynamics of the Ethereum staking market, but may fall short when considering the long-run perspective. We argue that the results presented above primarily reflect a short-term perspective on how the staking market might respond to changes in the issuance schedule. In the long run, the relative profitability of different staking solutions is likely to influence market shares, as new entrants or existing market participants opt for profitable staking methods and existing stakers exit unprofitable ones.
6.1 Short-term Perspective
@@
As argued above, the empirical estimates can be argued to be broadly consistent with the outcomes of the calibrated game-theoretic model, where solo stakers are modeled to experience disutility from variability in MEV revenues and where a portion of individuals staking through centralized providers are assumed to be inattentive. Under these assumptions, the staking supply of solo stakers is shown to adjust more strongly to changes in the issuance schedule compared to other types of stakers. Consequently, under the proposed issuance schedule, the market share of solo staking would be expected to decline, while the market share of centralized exchanges is shown to increase.
@@
This finding has important implications for the broader discussion around modifying the Ethereum issuance schedule. On the one hand, it suggests that a reduction in issuance could help limit the market share of liquid staking providers and the resulting dominance of LSTs within the Ethereum ecosystem. On the other hand, a reduction in the issuance schedule may come at the cost of an increased market share for centralized providers. In combination with a declining market share of solo stakers, this could contribute to a greater degree of centralization in the Ethereum validator set.
Finally, the results underscore that even with cost structures that would otherwise make solo stakers inelastic to changes in consensus yields, the competitive environment in the broader staking market can still make solo stakers vulnerable to being crowded out of the staking market when consensus yields fall. This is because even in a segregated staking market where existing solo stakers do not switch to other staking solutions at all, other market participants may be relatively insensitive to changes in consensus issuance and thus put pressure on solo stakers to reduce their staking supply when consensus yields fall.
This result also has important implications for the broader economic policy mix of the Ethereum protocol. It suggests that policies that help improve the competitive standing of solo stakers relative to other staking methods, for example the implementation of MEV burn, could help mitigate the otherwise negative effects of an issuance reduction on the market share of solo stakers.
6.2 Long-term Perspective
@@
However, exits and market participants switching between staking methods as a result of diminished profitability are not the only factors that could lead to diminished solo staking and an increasing dominance of liquid staking solutions in the market. This is because, over time, the divergence in profits between the different staking methods leads to a redistributive effect between the different staking methods. If stakers are able to reinvest their profits from staking, and staking profits are comparatively low for solo staking, their market share will continue to diminish. To illustrate this effect, we use the Nash equilibria of the baseline model outlined in Table 3 as a starting point and chart the cumulative changes in staking levels and staking ratios over time, assuming stakers reinvest their profits (revenues minus costs) and do not reevaluate their optimal level of staking supply over time. The results are shown in Figure 8. Figure 8(a) depicts the cumulative percentage change in the total staking level relative to the starting point. Similarly, Figure 8(b) depicts the cumulative change in the share of total ETH supply being staked. As such, Figure 8(b) also takes into account the effects of dilution.
Figure 8: Cumulative long-run effects of profit reinvestment
((a)) Staking level
((b)) Staking ratio
@@
7 Conclusion
Through empirical analysis and a game-theoretic model, we show that solo stakers are more responsive to changes in consensus issuance yields than ETH holders staking through centralized exchanges or liquid staking providers. As issuance declines, solo staking is expected to lose market share, while centralized exchanges gain dominance.
This heightened sensitivity of solo stakers persists despite their cost structures and is instead primarily driven by the competitive dynamics of the staking market. Other staking methods benefit from superior MEV access and DeFi yields, making them less responsive to issuance changes and creating a crowding-out effect on solo stakers. These findings underscore the importance of mechanisms such as MEV burn in mitigating the negative impact of issuance reductions on solo stakers.
@@ 7 Conclusion
The analysis presented in this paper has limitations. In particular, ike the assumptions used in calibrating the model. While the assumptions used in calibrating the theoretical model align with existing literature, they remain open to debate. Future research should prioritize refining the estimation of cost structures across different staking methods, which could enhance the accuracy and robustness of the model framework. Another important set of assumptions is that stakers are homogeneous within each type and that the ETH price remains constant over time (it is not affected by inflation and/or by the liquidity level). Relaxing these assumptions represents a promising direction for future research.
@@
We expect that the model developed here serves as a foundation for further research and extensions, providing a flexible framework for analyzing staking dynamics under various protocol adjustments. For tractability and compatibility with the broader model, we introduced several simplifying assumptions in the model extension with endogenous dSSP fees. However, we believe that our model could be expanded to facilitate a deeper exploration of the competitive dynamics between staking intermediaries.
Appendix A Derivation of Theoretical Model
Baseline model:
We solve for the equilibrium staking supply in the baseline model under the current issuance schedule ( the same approach is used to solve the extended models). In this case, the yield curve is:
$y_{i}(D_{T})=\frac{cF}{\sqrt{D_{T}}}=\frac{2.6\cdot 64}{\sqrt{D_{T}}}$
Each staker decides how much to stake taking into account other stakers’ behavior. Thus, we solve equations (3)-(1), getting
$\frac{\partial y_{i}(D)}{\partial\hat{d}_{ss}}$ $\cdot\hat{d}_{ss}+y_{i}(D)+y_{v}\cdot\frac{\partial P_{ss}(D)}{\partial\hat{d}_{ss}}-c_{ss}\cdot\frac{\partial\hat{d}_{ss}^{\alpha_{ss}}}{\partial\hat{d}_{ss}}=0,$
$(1-f_{d})$ $\cdot\left[\frac{\partial y_{i}(D)}{\partial\hat{d}_{t}}\cdot\hat{d}_{t}+y_{i}(D)+y_{v}\cdot\frac{\partial P_{t}(D)}{\partial\hat{d}_{t}}\right]+y_{d}-c_{t}\cdot\frac{\partial\hat{d}_{t}^{\alpha_{t}}}{\partial\hat{d}_{t}}=0,$
$(1-f_{c})$ $\cdot\left[\frac{\partial y_{i}(D)}{\partial\hat{d}_{r}}\cdot\hat{d}_{r}+y_{i}(D)+y_{v}\cdot\frac{\partial P_{r}(D)}{\partial\hat{d}_{r}}\right]-c_{r}\cdot\frac{\partial\hat{d}_{r}^{\alpha_{r}}}{\partial\hat{d}_{r}}=0,$
where
$y_{i}(D)$ $=\frac{2.6\cdot 64}{\sqrt{\hat{d}_{\theta}+(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta}}},$
$\frac{\partial y_{i}(D)}{\partial\hat{d}_{\theta}}\cdot\hat{d}_{\theta}$ $=-\frac{1}{2}\cdot\frac{2.6\cdot 64}{(\hat{d}_{\theta}+(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta})^{(3/2)}}\cdot\hat{d}_{\theta},$
$y_{v}\cdot\frac{\partial P_{\theta}(D)}{\partial\hat{d}_{\theta}}$ $=y_{v}\cdot\frac{(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta}}{(\hat{d}_{\theta}+(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta})^{2}},$
$c_{\theta}\cdot\frac{\partial\hat{d}_{\theta}^{\alpha_{\theta}}}{\partial\hat{d}_{\theta}}$ $=c_{\theta}\cdot\alpha_{\theta}\cdot\hat{d}_{\theta}^{\alpha_{\theta}-1}.$
The first three terms converge to zero for any $\hat{d}_{\theta}<<(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta}$ when $(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta}$ growth. However, $\frac{\partial y_{i}(D)}{\partial\hat{d}_{\theta}}\cdot\hat{d}_{\theta}$ converges faster than $y_{i}(D)$ and $y_{v}\cdot\frac{\partial P_{\theta}(D)}{\partial\hat{d}_{\theta}}$. Additionally,
$\frac{(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta}}{(\hat{d}_{\theta}+(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta})^{2}}\approx\frac{1}{(\hat{d}_{\theta}+(N_{\theta}-1)\cdot d_{\theta}+\sum_{-\theta}D_{-\theta})}=\frac{P_{\theta}(D)}{\hat{d}_{\theta}}.$
These points are exemplified in the following figures:
Figure 9: Illustration of Simplifying Assumptions
Note: x-Axis: Staking levels. y-Axis: $y_{i}(D)$ (blue ), $y_{v}\cdot P_{\theta}^{\prime}(D)$ (green), and $y_{i}^{\prime}(D)\cdot\hat{d}_{\theta}$ (orange). The figure above has $\hat{d}_{\theta}=100K$, while below $\hat{d}_{\theta}=32$.
Figure 10: Illustration of Simplifying Assumptions
Note: x-Axis: Staking levels. y-Axis: $y_{v}\cdot P_{\theta}^{\prime}(D)$ (blue) and $y_{v}\cdot P_{\theta}(D)/\hat{d}_{\theta}$ (orange). The figure above uses $\hat{d}_{\theta}=100K$, and below $\hat{d}_{\theta}=1^{\prime}000K$.
The previous observation allows us to simplify the FOCs and using $\hat{d}_{\theta}=d_{\theta}$ for each type in equilibrium, then the system of equations is, which can be solved numerically:
$y_{i}(D)+y_{v}\cdot\frac{P_{ss}(D)}{d_{ss}}=c_{ss}\cdot\frac{\partial d_{ss}^{\alpha_{ss}}}{\partial d_{ss}},\quad\text{for each ss}$
$(1-f_{d})\cdot\left[y_{i}(D)+y_{v}\cdot\frac{P_{t}(D)}{d_{t}}\right]+y_{d}=c_{t}\cdot\frac{\partial d_{t}^{\alpha_{t}}}{\partial d_{t}},\quad\text{for each t}$
$(1-f_{c})\cdot\left[y_{i}(D)+y_{v}\cdot\frac{P_{r}(D)}{d_{r}}\right]=c_{r}\cdot\frac{\partial d_{r}^{\alpha_{r}}}{\partial d_{r}},\quad\text{for each r}.$
Solos Stakers Expected MEV Revenues:
To model the variance over MEV rewards for moderate risk-averse solo-stakers, we can make
$\max_{\hat{d}_{ss}}\quad y_{i}(D)\cdot\hat{d}_{ss}+[y_{v}\cdot P_{ss}(D)-\frac{1}{2}\cdot y_{v}^{2}\cdot P_{ss}(D)\cdot(1-P_{ss}(D))]-(C_{ss}+c_{ss}\cdot\hat{d}_{ss}^{\alpha_{ss}}),$ (15)
where we are using
$E[Y_{MEV}]-Var[Y_{MEV}]$ $=y_{v}\cdot P_{ss}(D)-\frac{1}{2}\cdot y_{v}^{2}\cdot P_{ss}(D)\cdot(1-P_{ss}(D)).$
Notice that, since $P_{ss}(D)=\frac{d_{ss}}{D}\longrightarrow 0$ for small $d_{ss}$, then $E[Y_{MEV}]-Var[Y_{MEV}]\approx 0$. It follows that when we account for the higher variance in MEV rewards faced by solo stakers, a moderately risk-averse and small Solo Staker is assumed to exclude these revenues when deciding how much to stake, solving
$\max_{\hat{d}_{ss}}\quad y_{i}(D)\cdot\hat{d}_{ss}-(C_{ss}+c_{ss}\cdot\hat{d}_{ss}^{\alpha_{ss}}).$ (16)
Rather, we consider in the following a less risk-averse solo staker, with
$E[Y_{MEV}]-Var[Y_{MEV}]=y_{v}\cdot P_{ss}(D)-y_{v}^{2}\cdot P_{ss}(D)^{2},$
and for the first-order condition, we take the following approximation:
$\frac{\partial[y_{v}\cdot P_{ss}(D)-y_{v}^{2}\cdot P_{ss}(D)^{2}]}{\partial d_{ss}}\approx y_{v}\cdot\frac{P_{ss}(D)}{d_{ss}}\cdot(1-2\cdot y_{v}\cdot P_{ss}(D))$